OpenStudy (anonymous):

Just wondering about the concepts shouldn't you be able to use greens theorem to derive the polar conversion for double integrals?

OpenStudy (anonymous):

\[\int\limits_{\alpha}^{\beta}\int\limits_{a}^{b}f(rcos \theta, r \sin \theta)r dr d \theta\]

OpenStudy (anonymous):

why wouldn't you be able to create this from green's theorem using x=rcos0 and y =rsin0

OpenStudy (anonymous):

You should probably ask this question in maths group